Week of 060327

Open Actions from last week: Fill up c2sc4 staging disks

Chair: H.Renshall

On Call: Roberto

Monday:

Log: Major overheating problems in 513. Batch and CASTOR shut down. No FTS.

New Actions: Announce problem to GOCDB and via EDG broadcast.

Discussion: May last till Wednesday.

Tuesday:

Log:

New Actions:

Discussion:

Wednesday

Log:

Actions:

Discussion:

Thursday

Log:

Actions:

Discussion:

Friday

Log:

Actions:

Discussion:

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2006-03-28 - HarryRenshall
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback